if square root of sinx +cosx =0, then sin x=? Can u pls give me the solution as soon as possible?

Grade:11

2 Answers

Harshal Gupta
7 years ago
X = 3π/4You see, sin3π/4 + cos3π/4 = 1/√2 + -1/√2I guess you don`t remember all values, right?Anyways, try recalling most of them while practicing, it really helps
Soumendu Majumdar
6 years ago
root(sin x) + cos x = 0
implies sin x = cos^2 x …........by squaring both sides
implies sin^2 x + sin x – 1 = 0.........by using cos^2 x = 1 – sin^2 x
Now using shreedharacharya formula to find roots of a quadratic equation
you get
sin x = ( -1 ± root(5) )/2
now sin x= (-1 – root(5) )/2  is less than – 1 but sin x always lies within – 1 and 1
so  sin x = ( root(5) – 1 )/2
 
Hope it helps....